OpenStudy (anonymous):

You invest an initial $4,500 in an account that has an annual interest rate of 4.5%, compounded daily. How much money will you have in the account after 10 years? Round your answer to the nearest whole number.

OpenStudy (anonymous):

Hello! Are you familiar with the formula for compound interest?

OpenStudy (anonymous):

\[A = P(1 + \frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

if memory serves me correctly. @cebroskithe answer will be 7057.21

OpenStudy (anonymous):

@cebroski

OpenStudy (anonymous):

The principal, P, is the initial investment. P = 4500. The interest rate is r = 4.5% = 0.045. Compounded daily means n = 365 (ignoring leap year). And ten years means t = 10. Plug this into the equation and we get ...

OpenStudy (anonymous):

\[A = 4500(1 + \frac{ 0.045 }{ 365 })^{(365)(10)}\]

OpenStudy (anonymous):

Do you have any questions so far?